OpenStudy (anonymous):

3y – 16 = 14 and y is equal to one of the following: 9, 10, 11, 12, or 13. What is the value of y?

OpenStudy (anonymous):

10...

OpenStudy (perl):

Why guess if you can use algebra. 3y - 16 = 14 3y = 14 + 16 3y = 30 y = 10